# An Information Theoretic Perspective on Conformal Prediction

### v1

@@ Appendix C Proofs of Main Theoretical Results
In this section, we provide the proofs to our main results, namely the DPI bound in Proposition 3.1, the model-based Fano bound in Proposition 3.2, and the simple Fano bound in Corollary 3.1. For notational convenience, we use the shorthand $\alpha_{n}=\alpha-\nicefrac{{1}}{{1+n}}$ in most of the steps of the derivations.
C.1 DPI Bound
@@
We start with the DPI bound which we restate and proof below using the data processing inequalities discussed in Appendix B. Note that, when clear from the context, we remove explicit dependence on the calibration set ${\mathcal{D}}_{cal}$ from the derivations. It is implicitly assumed that the probability of the event $Y\in{\mathcal{C}}(x)$ is computed by marginalizing over ${\mathcal{D}}_{cal}$.
@@
Consider an arbitrary distribution $Q_{Y|X}$. Then we can use $P_{XY}$, and $P_{X}\times Q_{Y|X}$ in the data processing inequality for KL-divergence (Theorem B.4) to get:
@@
Remark C.2.
The derivation of the DPI bound places no assumptions on the conditional distribution $Q_{Y|X}$. However, in practice the underlying model $f$ often already provides such a distribution, and since it is typically trained to approximate $P_{Y|X}$ well, it makes sense to take $Q_{Y|X}$ as the distribution defined by the model itself. That is how we evaluate the DPI in all of our experiments. Finally, we can again use $H(Y|\hat{Y})$ instead of $H(Y|X)$ if the conformal method uses merely $\hat{Y}$.
Remark C.3.
@@
Remark C.4.
One of the appeals of the DPI bound is that the terms can be computed in a data-driven way using samples. While the cross-entropy can be estimated in an unbiased way with samples from the true distribution $P_{XY}$, we must be careful when estimating $Q(Y\in{\mathcal{C}}(x))$. The main challenge is that $Q(Y\in{\mathcal{C}}(x))$ appears inside a $\log$, and thus an empirical estimate $\hat{Q}(Y\in{\mathcal{C}}(x))$ would yield a lower bound of the negative KL divergence and would be biased. We can get an upper confidence bound on this estimate via the empirical Bernstein inequality Maurer & Pontil (2009), which we restate below.
Theorem C.5 (Empirical Bernstein Inequality Maurer & Pontil (2009)).
@@
C.2 Model-Based Fano Bound
@@
Remark C.6.
When the conformal prediction prediction is merely based on the model output $\hat{Y}=f(X)$, the model-based Fano bound can be modified to
@@
C.3 Simple Fano
In the derivation of the model-based Fano bound above, we placed no assumptions on the distribution $Q_{Y|X}.$ One simple choice that we consider in this section is the uniform distribution $Q_{Y|X}=\nicefrac{{1}}{{|{\mathcal{Y}}|}}$, akin to the classical Fano’s inequality Fano (1949) and the list decoding result in Proposition B.7.
@@ C.3 Simple Fano
Note that if we make an error ($E=0$) the correct class will not be inside ${\mathcal{C}}(X)$ and since $Q_{Y|X}$ is uniform the probability will be spread equally among the remaining $|{\mathcal{Y}}|-|{\mathcal{C}}(X)|$ labels, and we have $Q_{Y|X,{\mathcal{C}}(x),E=0}=\frac{1}{|{\mathcal{Y}}|-|{\mathcal{C}}(X)|}$. Through the same logic, we get that $Q_{Y|X,{\mathcal{C}}(x),E=1}=\frac{1}{|{\mathcal{C}}(X)|},$ and plugging both into (C.2) we get the simple Fano bound ∎
Remark C.7.
@@
Appendix D Further Theoretical Results and Proofs on the Prediction Set Size
D.1 Fano’s inequality for maximal prediction set size
@@
Now, we can plug both (22) and (D.1) into the data processing inequality of Theorem B.5. Rearranging the terms and getting the expectation from both sides w.r.t. ${\mathcal{D}}_{cal}$ and $X$ would yield:
@@
The first one is the Fano’s inequality for list decoding from Proposition B.7 while the second one yields the theorem. Note that if ${\mathcal{C}}(x)$ is an empty set, the probability $P(Y\in{\mathcal{C}}(x)|{\mathcal{D}}_{cal},X)$ is zero, and the term $P(Y\in{\mathcal{C}}(x)|{\mathcal{D}}_{cal},X)\log{|{\mathcal{C}}(x)|}$ disappears. Therefore, if we use the inequality $P(Y\in{\mathcal{C}}(x)|\sup_{{\mathcal{D}}_{cal}},X)\leq 1$, we need to introduce the term $[\log{|{\mathcal{C}}(x)|}]^{+}$ to keep the expectation well defined. ∎
D.2 Fano’s inequality for lower bound on prediction set size
@@
where $\hat{Q}^{0}_{Y|X}=q(y|x)\mathbb{I}[y\notin{\mathcal{C}}(x)]$ and $\hat{Q}^{1}_{Y|X}=q(y|x)\mathbb{I}[y\in{\mathcal{C}}(x)]$. When the conformal prediction prediction is merely based on the model output $\hat{Y}=f(X)$, the inequality can be modified to
@@
Remark D.3.
If we use the uniform distribution in the above bound, we get a bound similar to what is obtained from Fano’s inequality given in Proposition B.7, but with an additional factor of $\frac{1-\alpha}{1-\alpha+\frac{1}{n+1}}$. Since the factor is smaller than one, the current bound with the choice of uniform distribution is looser than Fano’s bound, although the gap vanishes for large $n$.
Appendix E Conformal Training
Split conformal prediction (SCP) Papadopoulos et al. (2002) has quickly become a popular framework for uncertainty estimation, largely thanks to its computational efficiency. One only needs access to a separate calibration data set to derive prediction sets with valid marginal coverage from any pretrained model. Given that training new machine learning models is becoming ever more time-consuming and expensive with new, larger architectures, this ability of applying conformal prediction to existing models is invaluable in a number of applications. Yet, it is reasonable to expect that the performance of the final set predictor could be improved if the conformal prediction process were to be accounted for during training of the model as well, steering the model towards better predictive efficiency, as it were. That is the motivation behind the line of work that we broadly refer to as conformal training Bellotti (2021); Stutz et al. (2022). In a nutshell, conformal training introduces a differentiable, and hence approximate, conformal prediction step during training so that one can directly optimize for desired properties of the set predictor, most notably its predictive efficiency.
To that best of our knowledge, Bellotti (2021) were the first to advance the idea of conformal training. The central step of their approach is to relax the prediction set defined by the model $f$ and define “soft” prediction sets $\hat{\mathcal{C}}_{f}(x)$, which contain each of the labels $y\in{\mathcal{Y}}$ with certain probability. That is, if ${\mathcal{C}}_{f}(x,y)\in\{0,1\}$ is the hard assignment of label $y$ to ${\mathcal{C}}(x)$, a corresponding soft version of this assignment can be defined as
@@ Appendix E Conformal Training
where $s_{f}(x,y)$ is the non-conformity score function defined by model $f$ evaluated at $(x,y)$, $\hat{q}$ is a thresholding value, $\sigma$ is the logistic sigmoid function, and $T$ is a temperature parameter controlling the smoothness of the soft assignment. Then we can define $\hat{\mathcal{C}}_{f}(x)$ as the vector collection all the soft assignments $\hat{\mathcal{C}}_{f}(x,y)$ for all labels $y\in{\mathcal{Y}}$. Similarly the size of $\hat{\mathcal{C}}_{f}(x)$ can be naturally defined as
@@
Bellotti (2021) then goes on to propose the following loss functions which are computed for each training batch ${\mathcal{B}}$
@@
Importantly, Bellotti (2021) argues the choice of the threshold $\hat{q}$ in (28) is immaterial since the model can learn to shift its logit values accordingly to match the constraints in ${\mathcal{L}}_{coverage}(f)$. We can then directly optimize (29) via stochastic gradient descent methods during training since it is fully differentiable with respect to the model parameters.
Stutz et al. (2022) build on the work of Bellotti (2021) by noticing that the calibration step is an important component in conformal prediction that should also be accounted for during training. To that end, they propose to split each training batch ${\mathcal{B}}$ in two: the ${\mathcal{B}}_{cal}$ half used for calibration, and the ${\mathcal{B}}_{test}$ used for testing. Now, instead of using an arbitrary threshold $\hat{q}$, we compute it using the quantile of ${\mathcal{B}}_{cal}$, or concretely
@@
With this modification, we no longer need to enforce valid coverage via ${\mathcal{L}}_{coverage}(f)$ and can optimize for low inefficiency directly by minimizing ${\mathcal{L}}_{size}(f)$ on ${\mathcal{B}}_{test}$. In that case, however, we only get a learning signal from ${\mathcal{B}}_{test}$, since the quantile operation applied to ${\mathcal{B}}_{cal}$ is non-differentiable. Stutz et al. (2022) bypass that limitation via differentiable sorting operators Cuturi et al. (2019); Blondel et al. (2020); Petersen et al. (2022), in particular via a version of differentiable sorting networks. In our experiments, we considered both fast sort Blondel et al. (2020) and the monotonic differentiable sorting networks of Petersen et al. (2022) but finally chose the latter since they proved more stable and provided richer gradient signals.
This version of their approach, which we refer to as ConfTr, only optimizes the size loss, but Stutz et al. (2022) also propose another variant which includes a classification loss term as follows
@@
Our own approach to conformal training follows the same recipe from Stutz et al. (2022), i.e., we also simulate a split conformal prediction step during training by splitting each training batch into two and using differentiable sorting operators. The key difference is in how we define the training objectives, which we derive from first principles and standard information theory inequalities. Not only do our upper bounds, DPI (4), MB Fano (5) and Fano (6), outperform the ConfTr objectives in a number cases, but they also do away with a few hyperparameters. Namely, the function $g$ in the ${\mathcal{L}}_{size}$, and hyperparameters controlling the relative importance of ${\mathcal{L}}_{size}$ and ${\mathcal{L}}_{class}$.
E.1 Deriving Conformal Training from Fano’s bound
Through Proposition 3.2, we can connect Fano’s bound for list decoding to the size loss from Stutz et al. (2022), proposed for conformal training. More specifically, by assuming a uniform distribution for $Q$ we can show that
@@ E.1 Deriving Conformal Training from Fano’s bound
Appendix F Experiments
In this section, we present further experimental results for conformal training in the centralized and federated setting. We start by defining the splits and architectures used for each data set, which are listed in Table 5. In most aspects, we follow the experimental design of Stutz et al. (2022), which serve as the main baseline against which we compare our models. All experimental results were obtained through our own implementation in Python 3 and Pytorch Paszke et al. (2017), which will be made publicly available upon acceptance for the sake of reproducibility.
@@ Appendix F Experiments
MNIST LeCun et al. (1998) 55K 5K 10K 50 1-layer MLP
Fashion-MNIST Xiao et al. (2017) 55K 5K 10K 150 2-layer MLP
EMNIST Cohen et al. (2017) 628K 70K 116K 75 2-layer MLP
CIFAR10 Krizhevsky et al. (2009) 45K 5K 10K 150 ResNet-34
CIFAR100 Krizhevsky et al. (2009) 45K 5K 10K 150 ResNet-50
Regarding the architectures, we also closely follow the experimental setup in Stutz et al. (2022). For MNIST we have a simple linear model, whereas for Fashion-MNIST and EMNIST we use 2-layer MLPs with 64 and 128 hidden units for first and second layers, respectively. For the CIFAR data sets, we use the default ResNet implementations from torchvision maintainers & contributors (2016), but changing the first convolution to have a kernel size of 3 and unitary stride and padding. We use Pytorch’s default weight initialization strategy for all architectures. For all datasets, we use a regular SGD optimizer with momentum $0.9$ and Nesterov gradients, accompanied by a step scheduler multiplying the initial learning rate by $0.1$ after $2/5$, $3/5$ and $4/5$ of the total number of epochs. We only use data augmentations on the CIFAR datasets, and differently from Stutz et al. (2022), we only apply random flipping and cropping for both CIFAR10 and CIFAR100.
F.1 Centralized Setting
We followed the experimental procedure of Stutz et al. (2022), and for each dataset and each method, we ran a grid search over the following hyperparameters using ray tune Moritz et al. (2018):
@@ F.1 Centralized Setting
  • Steepness of the differentiable sorting algorithm (monotonic sorting networks with Cauchy distribution Petersen et al. (2022)), which regulates the smoothness of the sorting operator; the higher the steepness value, the closer the differentiable sorting operator is to standard sorting. We considered steepness values in $\{1,10,100\}$.

### v2

@@ Appendix D Proofs of Main Theoretical Results
In this section, we provide the proofs to our main results, namely the DPI bound in Proposition 3.1, the model-based Fano bound in Proposition 3.2, and the simple Fano bound in Corollary 3.1. For notational convenience, we use the shorthand $\alpha_{n}=\alpha-\nicefrac{{1}}{{n+1}}$ in most of the steps of the derivations.
D.1 DPI Bound
@@
We start with the DPI bound which we restate and proof below using the data processing inequalities discussed in Appendix C. Note that, when clear from the context, we remove explicit dependence on the calibration set ${\mathcal{D}}_{cal}$ from the derivations. It is implicitly assumed that the probability of the event $Y\in{\mathcal{C}}(x)$ is computed by marginalizing over ${\mathcal{D}}_{cal}$.
@@
Consider an arbitrary distribution $Q_{Y|X}$. Then we can use $P_{XY}$, and $P_{X}\times Q_{Y|X}$ in the data processing inequality for KL-divergence (Theorem C.4) to get:
@@
Remark D.2.
The derivation of the DPI bound places no assumptions on the conditional distribution $Q_{Y|X}$. However, in practice, the underlying model $f$ often already provides such a distribution, and since it is typically trained to approximate $P_{Y|X}$ well, it makes sense to take $Q_{Y|X}$ as the distribution defined by the model itself. That is how we evaluate the DPI in all of our experiments. Finally, we can again use $H(Y|\hat{Y})$ instead of $H(Y|X)$ if the conformal method uses merely $\hat{Y}$.
Remark D.3.
@@
Remark D.4.
One of the appeals of the DPI bound is that the terms can be computed in a data-driven way using samples. While the cross-entropy can be estimated in an unbiased way with samples from the true distribution $P_{XY}$, we must be careful when estimating $Q(Y\in{\mathcal{C}}(x))$. The main challenge is that $Q(Y\in{\mathcal{C}}(x))$ appears inside a $\log$, and thus an empirical estimate $\hat{Q}(Y\in{\mathcal{C}}(x))$ would yield a lower bound of the negative KL divergence and would be biased. We can get an upper confidence bound on this estimate via the empirical Bernstein inequality [39], which we restate below.
Theorem D.5 (Empirical Bernstein Inequality [39]).
@@
D.2 Model-Based Fano Bound
@@
Remark D.6.
When the conformal prediction is merely based on the model output $\hat{Y}=f(X)$, the model-based Fano bound can be modified to
@@
D.3 Simple Fano
In the derivation of the model-based Fano bound above, we placed no assumptions on the distribution $Q_{Y|X}.$ One simple choice that we consider in this section is the uniform distribution $Q_{Y|X}=\nicefrac{{1}}{{|{\mathcal{Y}}|}}$, akin to the classical Fano’s inequality [18] and the list decoding result in Proposition C.7.
@@ D.3 Simple Fano
Note that if we make an error ($E=0$) the correct class will not be inside ${\mathcal{C}}(X)$ and since $Q_{Y|X}$ is uniform the probability will be spread equally among the remaining $|{\mathcal{Y}}|-|{\mathcal{C}}(X)|$ labels, and we have $Q_{Y|X,{\mathcal{C}}(x),E=0}=\frac{1}{|{\mathcal{Y}}|-|{\mathcal{C}}(X)|}$. Through the same logic, we get that $Q_{Y|X,{\mathcal{C}}(x),E=1}=\frac{1}{|{\mathcal{C}}(X)|},$ and plugging both into (D.2) we get the simple Fano bound ∎
Remark D.7.
@@
Appendix E Further Theoretical Results and Proofs on the Prediction Set Size
E.1 Fano’s inequality for maximal prediction set size
@@
Now, we can plug both (23) and (E.1) into the data processing inequality of Theorem C.5. Rearranging the terms and getting the expectation from both sides w.r.t. ${\mathcal{D}}_{cal}$ and $X$ would yield:
@@
The first one is the Fano’s inequality for list decoding from Proposition C.7 while the second one yields the theorem. Note that if ${\mathcal{C}}(x)$ is an empty set, the probability $P(Y\in{\mathcal{C}}(x)|{\mathcal{D}}_{cal},X)$ is zero, and the term $P(Y\in{\mathcal{C}}(x)|{\mathcal{D}}_{cal},X)\log{|{\mathcal{C}}(x)|}$ disappears. Therefore, if we use the inequality $P(Y\in{\mathcal{C}}(x)|\sup_{{\mathcal{D}}_{cal}},X)\leq 1$, we need to introduce the term $[\log{|{\mathcal{C}}(x)|}]^{+}$ to keep the expectation well defined. ∎
E.2 Fano’s inequality for lower bound on prediction set size
@@
where $\hat{Q}^{0}_{Y|X}=q(y|x)\mathbb{I}[y\notin{\mathcal{C}}(x)]$ and $\hat{Q}^{1}_{Y|X}=q(y|x)\mathbb{I}[y\in{\mathcal{C}}(x)]$. When the conformal prediction is merely based on the model output $\hat{Y}=f(X)$, the inequality can be modified to
@@
Remark E.3.
If we use the uniform distribution in the above bound, we get a bound similar to what is obtained from Fano’s inequality given in Proposition C.7, but with an additional factor of $\frac{1-\alpha}{1-\alpha+\frac{1}{n+1}}$. Since the factor is smaller than one, the current bound with the choice of uniform distribution is looser than Fano’s bound, although the gap vanishes for large $n$.
Appendix F Conformal Training
Split conformal prediction (SCP) [45] has quickly become a popular framework for uncertainty estimation, largely thanks to its computational efficiency. One only needs access to a separate calibration data set to derive prediction sets with valid marginal coverage from any pretrained model. Given that training new machine learning models is becoming ever more time-consuming and expensive with new, larger architectures, this ability to apply conformal prediction to existing models is invaluable in a number of applications. Yet, it is reasonable to expect that the performance of the final set predictor could be improved if the conformal prediction process were to be accounted for during training of the model as well, steering the model towards better predictive efficiency, as it were. That is the motivation behind the line of work that we broadly refer to as conformal training [8, 59]. In a nutshell, conformal training introduces a differentiable, and hence approximate, conformal prediction step during training so that one can directly optimize for desired properties of the set predictor, most notably its predictive efficiency.
To the best of our knowledge, Bellotti [8] were the first to advance the idea of conformal training. The central step of their approach is to relax the prediction set defined by the model $f$ and define “soft” prediction sets $\hat{\mathcal{C}}_{f}(x)$, which contain each of the labels $y\in{\mathcal{Y}}$ with a certain probability. That is, if ${\mathcal{C}}_{f}(x,y)\in\{0,1\}$ is the hard assignment of label $y$ to ${\mathcal{C}}(x)$, a corresponding soft version of this assignment can be defined as
@@ Appendix F Conformal Training
where $s_{f}(x,y)$ is the non-conformity score function defined by model $f$ evaluated at $(x,y)$, $\hat{q}$ is a thresholding value, $\sigma$ is the logistic sigmoid function, and $T$ is a temperature parameter controlling the smoothness of the soft assignment. Then we can define $\hat{\mathcal{C}}_{f}(x)$ as the vector collection of all the soft assignments $\hat{\mathcal{C}}_{f}(x,y)$ for all labels $y\in{\mathcal{Y}}$. Similarly the size of $\hat{\mathcal{C}}_{f}(x)$ can be naturally defined as
@@
Bellotti [8] then goes on to propose the following loss functions which are computed for each training batch ${\mathcal{B}}$
@@
Importantly, Bellotti [8] argues the choice of the threshold $\hat{q}$ in (29) is immaterial since the model can learn to shift its logit values accordingly to match the constraints in ${\mathcal{L}}_{coverage}(f)$. We can then directly optimize (30) via stochastic gradient descent methods during training since it is fully differentiable with respect to the model parameters.
Stutz et al. [59] build on the work of Bellotti [8] by noticing that the calibration step is an important component in conformal prediction that should also be accounted for during training. To that end, they propose to split each training batch ${\mathcal{B}}$ in two: the ${\mathcal{B}}_{cal}$ half used for calibration, and the ${\mathcal{B}}_{test}$ used for testing. Now, instead of using an arbitrary threshold $\hat{q}$, we compute it using the quantile of ${\mathcal{B}}_{cal}$, or concretely
@@
With this modification, we no longer need to enforce valid coverage via ${\mathcal{L}}_{coverage}(f)$ and can optimize for low inefficiency directly by minimizing ${\mathcal{L}}_{size}(f)$ on ${\mathcal{B}}_{test}$. In that case, however, we only get a learning signal from ${\mathcal{B}}_{test}$, since the quantile operation applied to ${\mathcal{B}}_{cal}$ is non-differentiable. Stutz et al. [59] bypass that limitation via differentiable sorting operators [16, 9, 48], in particular via a version of differentiable sorting networks. In our experiments, we considered both fast sort [9] and the monotonic differentiable sorting networks of [48] but finally chose the latter since they proved more stable and provided richer gradient signals.
This version of their approach, which we refer to as ConfTr, only optimizes the size loss, but Stutz et al. [59] also proposes another variant which includes a classification loss term as follows
@@
Our own approach to conformal training follows the same recipe from [59], i.e., we also simulate a split conformal prediction step during training by splitting each training batch into two and using differentiable sorting operators. The key difference is in how we define the training objectives, which we derive from first principles and standard information theory inequalities. Not only do our upper bounds, DPI (4), MB Fano (5) and Fano (6), outperform the ConfTr objectives in a number cases, but they also do away with a few hyperparameters. Namely, the function $g$ in the ${\mathcal{L}}_{size}$, and hyperparameters controlling the relative importance of ${\mathcal{L}}_{size}$ and ${\mathcal{L}}_{class}$.
F.1 Deriving Conformal Training from Fano’s bound
Through Proposition 3.2, we can connect Fano’s bound for list decoding to the size loss from [59], proposed for conformal training. More specifically, by assuming a uniform distribution for $Q$ we can show that
@@ F.1 Deriving Conformal Training from Fano’s bound
Appendix G Experiments
In this section, we present further experimental results for conformal training in the centralized and federated setting. We start by defining the splits and architectures used for each data set, which are listed in Table 4. In most aspects, we follow the experimental design of [59]. All experiments were conducted on commercially available NVIDIA GPUs using our own implementation in Python 3 and Pytorch [47], which will be made publicly available upon acceptance for the sake of reproducibility. All data sets were retrieved directly from torchvision [38].
@@ Appendix G Experiments
MNIST [27] 55K 5K 10K 50 1-layer MLP
Fashion-MNIST [67] 55K 5K 10K 150 2-layer MLP
EMNIST [11] 628K 70K 116K 75 2-layer MLP
CIFAR10 [23] 45K 5K 10K 150 ResNet-34
CIFAR100 [23] 45K 5K 10K 150 ResNet-50
Regarding the architectures, we also closely follow the experimental setup in [59]. For MNIST we have a simple linear model, whereas for Fashion-MNIST and EMNIST we use 2-layer MLPs with 64 and 128 hidden units for first and second layers, respectively. For the CIFAR data sets, we use the default ResNet implementations from torchvision [38], but changing the first convolution to have a kernel size of 3 and unitary stride and padding. We use Pytorch’s default weight initialization strategy for all architectures. For all datasets, we use a regular SGD optimizer with momentum $0.9$ and Nesterov gradients, accompanied by a step scheduler multiplying the initial learning rate by $0.1$ after $2/5$, $3/5$ and $4/5$ of the total number of epochs. We only use data augmentations on the CIFAR datasets, and differently from [59], we only apply random flipping and cropping for both CIFAR10 and CIFAR100.
G.1 Centralized Setting
We followed the experimental procedure of [59], and for each dataset and each method, we ran a grid search over the following hyperparameters using ray tune [43]:
@@ G.1 Centralized Setting
  • Steepness of the differentiable sorting algorithm (monotonic sorting networks with Cauchy distribution [48]), which regulates the smoothness of the sorting operator; the higher the steepness value, the closer the differentiable sorting operator is to standard sorting. We considered steepness values in $\{1,10,100\}$.